OpenStudy (anonymous):

***WILL MEDAL*** A sequence is defined by the formula f(n + 1) = f(n) – 3. If f(4) = 22, what is f(1)?

OpenStudy (trojanpoem):

Let n = 3 f(3+1) = f(3) - 3 f(4) = f(3) - 3 f(3) = f(4) + 3 = 22 + 3 = 25 Let n =2 f(3) = f(2) - 3 f(2) = f(3) + 2 = 25 + 3 = 28 Let n = 1 f(2) = f(1) - 3 f(1) = f(2) + 3 = 28 + 3 = 31

OpenStudy (trojanpoem):

Another solution: f(n+1)- f(n) = -3 = d d = -3 (common difference) and we have a sequence which a_{4} = 22 (4th term) a_{4} = a_{0} + d(n-1) 22 = a_{0} -3(4-1) a{0} = 22 + 9 = 31